Left Direct Gregorian Technical Observing with IRTC: UT 2009 October 07

Observer: JHill (Tucson), DMiller (LBTO)
Telescope Operator: Aaron Ceranski
SW support: TLeibold, N"Something or another" (Tucson)


Closed all night due to clouds, high humidity and high winds.

Two hours were spend finding an unplugged USB cable on the back of the IRTC (see IssueTrak # 2381)


All Times in UT

00:30 Tried to communicate with IRTC but failed. The Bologna SW is not able to communicate to the camera. Dave T had John L power cycle the IRTC power/communication box and the IRTC computer earlier in the day and he was not able to talk to the IRTC either. I tried several power cycles to no avail.

01:00 Scattered clouds at sunset

1:15 Brian and Jeff checked one power cable at the IRTC power bar and found a voltage.

1:20 Aaron deploys the AO Anemometer on DX M2 swing arm and retracts LBC Red at Doug's request so wind measurements at the future position of the Adaptive secondary can be collected.

1:30 Doug and Jeff changed from fiber 1&2 to 5&6. The lights on the back of the IRTC power supply in the tree house change from green to red during the change of fiber and changed back to green when the new fibers were installed.

2:00 Doug tried another IRTC Power cycle and reboot. No Luck

2:15 Doug connects the USB port from the Xeva Camera directly to his laptop (Mac running VmWare with WIndows XP). This bypasses the USB extender. If the camera is on, the USB port should be recognized as Xenics Drives. No USB device is seen. Either the camera does no t have power or the USB cable is unplugged

2:30 John L and Brian check IRTC connections (for the fourth time today). All looks fine

3:00 Direct Gregorian rotator is not responding. Full power cycle of rotator electronic in upper DX treehouse solves the problem

3:15 John L and Aaron rebalance telescope

3:30 Execute presetTelescope to test all systems are functioning. Successful. The preset fails when GCS cannot find a star in a saturated image (as expectec). Norm and Torsten go home.

4:00 Aaron tips telescope to Horizon and Tim fills the LBC Dewars.

4:45 Doug check the Hardware entries on the IRTC computer to see if the IRTC USB connection is recognized. It is not.

4:30 Tim and Doug use scissor lift to examine the USB plug in the back of the Xeva Camera (IRTC). When Tim pushed on the USB cable it was clear that it had come unplugged (it pushed several inches into the housing and should not have moved at all). They were able to get the foam rubber cover opened enough to get the mini-USB plug back in. The only strain relief for the USB cable was that is was taped to the power cable. It looked like the tape had slipped a bit and allowed the USB cable to unplug. Plus, if the power cable was bent at 90 degrees, that would probably unplug the USB cable. I would guess the power to USB cable got bumped during dismounting or mounting and unplugged the USB. I don't think this is a handling problem because it would not take much to unplug the USB cable. I think this is a design/lack of proper strain relief problem. Tim and Doug used two tie wraps to secure the USB cable to the power cable so it will be a bit more difficult to unplug the USB, but this is not a robust solution. See IssueTrak # 2381

The attached image is the Hardware Devices on the IRTC computer (Settings =>Control Panel => System => Hardware => Device Manager). Note, the USB driver for the Xeva Camera is listed as Xenics Drivers => XEVA-CL-FPA-320-60Hz. The driver for the PCI fast data acquisition card is the Other Devices > CC Driver.

  • IRTC Hardware Device Manager Output with IRTC working properly:

I think if there is no power to the camera the Other Devices will be displayed in the Hardware device list, but no CC Driver, and the Xenics Drivers will not be shown, only USB Root Hubs. It the camera has power but the USB cable has come unplugged, the CC Driver will be displayed but the Xenics Drivers will not be shown. Instead, there will be an USB device listed with "Unknown Device". This is because the USB converter will still be powered and the IRTC computer will something plugged in, but not recognize it.

5:30 Doug and John H test the "irc OffsetGuiding rot ra dec RADEC right" command. IIF and PCS report they received and executed the request. They think a script can be used to send at 1-4 Hz to "simulate" non-sidereal tracking. Doug will write a quick TclTk script to try. When we get on the sky we will have to see if the offset is relative or absolute. We will have to test rate, amount and direction of the correction on sky.

6:00 Dave T sends an email with the estimated correct x-flip values. Doug changes the oacontrol.conf values and restarts the aocserver. He also modifies the GSC cfg file and Aaron restarts GCS.

11:00 FIRE ALARM!!! Found no fire. Only one sensor went off in the high bay and is now OK after a reset. Maybe water? Talk about some groggy puffy eyed boys.

11:10 Wet outside. Rain sensor says yes. Wind gusting to 20 m/s

12:04 Time to sleep

-- DougMiller - 29 Sep 2009
I Attachment Action Size Date Who Comment
Dave_x_flip_email.pdfpdf Dave_x_flip_email.pdf manage 30 K 07 Oct 2009 - 08:11 DougMiller Dave's X-Flip Email
IRTC_Drivers.jpgjpg IRTC_Drivers.jpg manage 138 K 07 Oct 2009 - 07:57 DougMiller IRTC Hardware Device Manager Output with IRTC working properly
daq.20091007.gifgif daq.20091007.gif manage 31 K 08 Oct 2009 - 03:54 JohnHill Plot of Telescope and Mirror Temperatures
Topic revision: r5 - 08 Oct 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback